Brandon is an analyst at a wealth management firm. One of his clients holds a $7,500 portfolio that consists of four stocks.According to Brandons recommendation, assuming that the market is in equilibrium, how much will the portfolios required retSuppose instead of replacing Atteric Inc.s stock with Transfer Fuels Co.s stock, Brandon considers replacing Atteric Inc.s
Answer #1

Change in Required Return = w(Change in Beta)(Market Risk Premium)

Change in Required Return = (0.35)(0.50 - 0.75)0.055

Change in Required Return = 0.4840%

New Required Return = 0.08829 - 0.004840 = 8.345%

Shares are overvalued as expected return rate(6.85%) is lower than Required Return(8.345%)

Increase.